FEL Upgrade Project Weekly Brief
May 7-11, 2001

We hosted a visit to the FEL, the cavity fabrication facility, and the optical klystron test area by Cdr. Roger McGinnis, Neal Barron, Quenton Salter, and two naval architects to determine the feasibility of adapting our design to a shipboard configuration.  There is a lot of work to do but no showstoppers were uncovered.

Planning for the June workshop continued including trying to optimize plans for scale up of the system and the necessary technology developments required.

Abstracts for the FEL conference were prepared and reviewed prior to submission by the deadline 0f 5/11/01.

WBS 4 (Injector):
Installed a 1um SST faraday shielded implanted electrode into the FET system. Setting the system up for bake prior to measurements. Continued detail efforts on the cathode support tube and stalk.  The first cut of chamber drawings are in review.

WBS 6 (RF):
Zone 1 - Operational, but was blowing the -18v fuse for position 3 at random times.  The RF Crate was replaced and now the RF Control Module does not communicate with CAMAC.  This is being addressed this afternoon by AES.

Zone 2 - Operational with no problems.

Zone 3 - No progress.  Waiting for additional parts.

Zone 4 - Complete, except circulator for position 8 is out for repair.  This zone is now being used to test klystron beam current stability by controlling the heater voltage.

Gun HVPS - Vendor is planning to ship the upgraded HVPS by the end of May.  The design of the HVPS Tank is in progress.  A new cylindrical tank is being considered as a replacement for the elongated tank design.  The design of a new elbow for the transmission line is just getting started.

Quarter HVPS - PSC has submitted a quote of $279K total to upgrade the 2 HVPS's for the Quarter. I am still contacting transformer vendors for quotes to see if we can do this task in-house.  A meeting is scheduled for next Monday to see if HPEE can fit this effort into their work schedule.

WBS 8 (I&C):
The old "Area 51" is being re-commissioned for "clean" beam diagnostics assembly.  The beam viewers will be setup with solenoids, limit switches and air lines prior to being sent to the Test Lab for girder assembly.  This will relieve the congestion in the tunnel next year during installation and be a much better utilization of manpower now and next year.  A clean hood and seal-a-meal was ordered and is being fitted on one of the benches for work on viewers.  A dry
N2 line was run from Lab 6 to fill the bags prior to sealing.
The Molectron quad detector was tested; it performed very well.  This will be used for optical transport control.  The CW test with the drive laser and a chopper was interrupted due to a water leak. This will be done next week.  The circuit design for "daughter board" sample & hold has been submitted to EECAD. This board will fit to the top of the quad preamp board inside the shielded enclosure.
EPICS software was written for a CNC-like script which a user can use to program the FEL to perform a fairly general set of drilling motions.  This has not yet been tested.  Also the Pulse Truncator software and IP pac was installed this will be tested beginning next week, independent of the controller hardware.
The last of the LSS work is being finished up.  New status signs were put on the doors of labs 5 and 6. The new hutch in lab 6 had limit switches put on the doors.  There are 12 doors, each has two limit switches - one for each corner opposite the hinge.
The HP workstation in lab 1 was moved to accommodate shelving for assorted laser experimental equipment.  This is being done to clean up the "piles" around the labs.  Thanks to all for helping in the cleanup.  The MCC workstation will be moved into Lab 5 to support VME board testing.  Parts continue to trickle in for the Ion Pump Power Supplies.  Assembly and Fabrication Drawings are nearly complete and Schematics will be completed this week.  I will be sending those out for assemble quotes next week as well.
I have continued to update drawings.  The check plots for the LSS Lab Interface Box update and the Charge/Dump Current Monitor Chassis were returned to EECAD today.  Documentation of the picomotor relay chassis is in progress.  The wiring diagram for the picomotor controller chassis has been signed off and forwarded to document control.  Drawings were completed for the Camera Power Supply Chassis for the user labs and those chassis panels have been sent out for machining and silkscreen. Deborah met with Jim Coleman concerning the Machine Protection System documentation and we will be diving into that next week.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
•  DULY Research continues adapting their magnetic model of the small injector Dipole (GV) into a practical
    hardware design.
• They continue to work the details and assembly of the large injector dipole (GU).
Optical Chicane Dipole (DW)
• The design package is finished backchecking for the second time.  I will give it a final look before signature.
•  Innovation Services reviewed our coil specification for dipole insulation, bringing it up to date for resin systems
    and enhancing the quality control provisions.  They are now working on a final report that will serve as an in
    house reference on how they arrived at their conclusions embodied in the revised specification.
Arc Dipoles (GY, GX, GQ)
• The material for the pole tips of the GY & GX was shipped.
•  The magnetic analysis of the GY stopped until Dave Douglas can consider the options implicated by the latest
    results (3 mm parallel off sets on each side) with path length correctors and synchrotron light periscopes.  Tom
    Schultheiss is planning a trip to JLab to finalize the magnetic analysis.
• We are checking the 3/3 inch stack of drawings of the GY magnet.
• They continuing design and magnetic modeling the Arc Bend Dipole (GX).  In the design area, the coils were
    modeled and the field clamps were developed.  The magnetic model is about finished so that we can start
QG (3 inch quad)
• The coil-manufacturing package remains in Procurement cycle.
• The core drawings are still in the final stages of polish as small points were brought up during a review this
• Our procurement strategy, as worked out with Procurement in a meeting this week, is to pre qualify vendors
    from our list of 8 with a “heads up” Request for Information (RFI) – with packages going out this week.  This
    will shorten the ultimate bid time and eliminate spurious bids. For the future Request for Proposal, instead of
    judging the proposals by an arduous “best value” method, we decided that the submittals required by the
    drawing and specification are simple enough to judge by a Pass/Fail process. Final selection of vendor is
    determined by price.
• During the week, the prototype quadrupole had two cycles of chamfer cutting to reduce the dodecupole, our
    final requirement of the prototype process.  The results indicate one more cut, to a .187 inch chamfer will meet
    David Douglas’s specification.
QX (Trim Quad)
• The 3D magnetic model was debugged and is now yielding good results.  Now detailed magnetic design can
    start, moving the protrusion of the coil ends in the horizontal or vertical to minimize the higher order multipoles.
• We received the conductor that was specially ordered for these coils.  It was slightly out of our expected tight
    tolerances but may now be sequestered and the design based on it, which will significantly reduce the variability
    of the final magnets.
Corrector Dipoles (DB & DJ)
• The package continues in vendor’s hands for a re-bid.
• The Dipole Steel continues in Procurement cycle.
• We continued design layout of the girders between cryomodule, anticipating a series of working sessions next
    week where all parties have the graphics available to help condense 5 lbs. into a 2 lb. bag.
• We are putting the Task definition together for design of the arc chambers.  The overall configuration of ports,
    thickness, intersections and diagnostics is clarified.
• The lower half of the new alignment cartridges were put on order.

WBS 10 (Wiggler):
Fabrication of the corrections coils for the wigglers has started.  The design of the coils is being incorporated into the overall design.  Most of the drawings for the wiggler measurement jig have been signed off.  Characterization of the dispersion section with a corrector coil continues.

WBS 11 (Optics):
The ultraviewer and near-Brewster window drawing packages are out for bid, we are already ordering associated optics for the assemblies.  The design focus shifts back to the optical cavity assembly internals, in particular the gimbal mounts.  We have generated more detailed specs on the linear motion needed to orient the mirrors, and have started detailing the linear actuation.  We anticipate the beam-path shields for the mirror test stand to arrive today (5/11).

Since CLEO (the Conference on Lasers and Electro-Optics) was in Baltimore, a group of us (including the entire optics group), spent a day at the exhibit show.

We lost a seal for the Drive Laser's Nd:YLF laser rod.  After clean up and cleaning of optics, a new rod will be installed.

Work continued to define the "killer" demonstration of our high flux X-ray source.  We believe a pump/probe experiment looking at resonant phonon decay in crystals might be the answer.  We are continuing to analyze the requirements.

We prepared hardware for installation Monday of   a) a diamond window to allow extraction of far IR/mm waves, b) a new X-ray diffraction crystal, and c) silicon viewers before and after the first 180 bend to allow careful CSR emittance growth measurements to be performed at a later date.